What is the Area and slope of: velocity -time graph distance vs time acceleration vs time graph pls answer sir

Grade:12

1 Answers

erra akhil
8 years ago
Dear Azx,
Slope implies differentiation of x and y axes.
So,in case 1:velocity vs time graph
Slope=dv/dt=instantaneous acceleration.
Area  under this graph gives total distance travelled.
case 2:distance vs time
slope=ds/dt=instantaneous velocity
Area under distance time graph is meaning less.
case 3:accelearion vs time
slope=da/dt=rate of change of acceleration
area= average velocity